# Data-Augmented Deep Learning for Downhole Depth Sensing and Field Validation

### v2

@@ Abstract
Accurate downhole depth measurement is essential for oil and gas well operations, directly influencing reservoir contact, production efficiency, and operational safety. Collar correlation using a casing collar locator (CCL) is fundamental for precise depth calibration. While neural network-based CCL signal recognition has achieved significant progress in collar identification, preprocessing methods for such applications remain underdeveloped. Moreover, the limited availability of real well data poses substantial challenges for training neural network models that require extensive datasets. This paper presents a system integrated into downhole tools for CCL signal acquisition to facilitate dataset construction. We propose comprehensive preprocessing methods for data augmentation and evaluate their effectiveness using our neural network models. Through systematic experimentation across various configuration combinations, we analyze the contribution of each augmentation method. Results demonstrate that standardization, label distribution smoothing (LDS), and random cropping are fundamental requirements for model training, while label smoothing regularization (LSR), time scaling, and multiple sampling significantly enhance model generalization capability. The F1 scores of our two benchmark models trained with the proposed augmentation methods maximumly improve from 0.937 and 0.952 to 1.0 and 1.0, respectively. Performance validation on real CCL waveforms confirms the effectiveness and practical applicability of our approach. This work addresses the gaps in data augmentation methodologies for training casing collar recognition models in CCL data-limited environments.
@@ I Introduction
Figure 1: Cross-sectional illustration of a typical oil and gas well structure. Representative casing collar signatures from magnetic response are illustrated in dark blue near the corresponding casing collar, while the typical interference signals distant from casing collars are illustrated in dark green. Refer to figures from [1].
Accurately positioning downhole tools – including perforating guns, bridge plugs, and packers – relative to target zones is essential in modern oil and gas well operations, directly affecting reservoir contact, production efficiency, and operational safety [2]. Central to this task is precise downhole depth measurement, a challenge compounded by the extreme geometries of wellbores, as shown in Fig. 1 organized from [1], which often span thousands of meters in length while maintaining diameters of only a few inches [3].
Although a length measuring wheel (LMW) is able to estimate depth as tools descend during wireline intervention operations, its accuracy remains insufficient. To improve depth measurement accuracy, calibration using signals from a casing collar locator (CCL) – a magnetic downhole positioning tool that detects magnetic anomalies at the casing collars of wellbores or pipes – is fundamental [4]. The CCL produces a characteristic magnetic response as an electrical signal when passing through each collar [4, 5, 6], termed a “collar signal” or “collar (magnetic) signature”. The characteristic magnetic response pattern typically exhibits a bimodal distribution, as illustrated by dark blue waveforms in Fig. 1. Through collar correlation, which refers to matching CCL logs with casing tally, depth reference markers are established, enabling actual and accurate depth measurement [7, 6]. The casing tally, also known as the list of collars or casing string reference depths, records the depths of casing collars and is commonly extracted from cementing quality data.
While collar correlation using CCL signals is an established method, collar signature recognition presents significant challenges. CCL signal integrity can be severely compromised by multiple factors, including cable effects, wellbore conditions [8], tool motion (swing or rotation), amplifier saturation, and environmental noise [7, 9]. Consequently, collar signature waveforms become increasingly ambiguous, necessitating robust recognition of collar signatures amid various interferences, as illustrated by dark green waveforms in Fig. 1.
@@
The limited sample amount and unique characteristics of CCL signal data necessitate specialized preprocessing methods for neural network training. However, research on preprocessing methods remain scarce, despite extensive work on collar signature recognition [4, 9, 11, 12, 13, 14, 17, 19, 18]. Fortunately, extensive work exists on data augmentation for preventing overfitting and improving generalizability when training on small datasets. Data normalization methods – including min-max scaling, Z-score normalization (standardization), and robust scaling – demonstrate important roles in stabilizing input distributions and gradients, preventing vanishing and exploding gradients, and ultimately accelerating training convergence [24, 25, 26, 27]. Label smoothing regularization (LSR) and label distribution smoothing (LDS) discourage overconfident predictions, thereby enhancing model generalizability [28, 29]. Similarly, employing probability maps for boundary prediction instead of one-hot encoding (OHE) labels through boundary probabilization transforms the training objective from single-point prediction to probability distribution estimation, enabling models learn smooth and fuzzy decision boundaries that are more resistant to noise or perturbations [28, 15, 30]. Notably, Gaussian kernel often deliver optimal results in LDS applications [28]. Furthermore, graph augmentation techniques including randomly cropping, scaling, translation, and noise injection expand training datasets while improving model generalizability and robustness [31, 32].
@@
  • We develop a system integrated into downhole tools, called the Signal Collecting Vessel (SCV), illustrated in Fig. 2. The SCV samples original CCL signal within wellbores or pipes, converts them to digital format, and stores them as waveforms for dataset construction.
  • We propose two neural networks models for collar signature recognition that serve as benchmarks for evaluating data preprocessing methods, as illustrated in Fig. 3. The first model, Thin AlexNet (TAN), is modified from AlexNet – a classic and proven architecture in pattern recognition. The second model, Miniaturized AlexNet (MAN), is a simplified version of TAN with fewer layers.
@@
  • We conduct extensive experiments across various configuration combinations to validate our methods. Results demonstrate that standardization, LDS, and random cropping are fundamental requirements for models training, while LSR, time scaling, and multiple sampling significantly enhance model generalization capability.
@@ II-A Problem Transformation
As previously discussed, accurately positioning downhole tools through collar correlation relies on correct identification of bimodal patterns (i.e., collar signatures) within CCL signals. The center of each bimodal pattern is typically considered as the moment when the CCL coincides with the casing collar.
By observing signal characteristics, collar signatures can be clearly identified from signals proximate to the corresponding collar rather than from distant signals. Therefore, collar signature recognition is largely independent of CCL signals on large time scales. Consequently, focusing on CCL signal fragments using an appropriately sized sliding window is reasonable, consistent with approaches in [15] and [16]. When CCL signals are sampled at a fixed frequency, the absolute value of timestamps within sliding windows becomes irrelevant, indicating that the pure CCL signal sequence sufficiently conveys the necessary information.
As shown in Fig. 1, the casing tally indicates a series of depths that can be mapped to collar signatures on the signal timeline as a series of marks. The center points of collar signatures marked on the timeline (hereafter referred to as “collar marks”) constitute a series of timestamps that require different representation when captured by sliding windows. The most direct method for representing collars marks is one-hot encoding (OHE), where 1 denotes “is a collar mark” and 0 denotes “is not a collar mark”. This approach transforms collar mark prediction into a binary classification problem.
@@
However, collar marks exhibit extreme sparsity. Consequently, the number of 0s exceeds that of 1s by several orders of magnitude, representing severe class imbalance. Due to the nature of backpropagation neural networks (BPNNs), gradients become spare and penalties are unevenly distributed, causing training instability and slow convergence. This ultimately renders effective classifier training nearly impossible. To address this issue, one-hot encoding is replaced with a probability map [33, 15], where labels represent the probability of boundary occurrence. Interestingly, according to the theory in [33], label represents the distribution of relative importance across possible categories (i.e., “is a collar mark”). This replacement transforms boundary classification into boundary membership classification, providing meaningful feedback during backpropagation and ultimately enabling stable and efficient network training.
II-B Collecting the original CCL waveforms
@@
Figure 2: Structure of downhole tools integrated with the Signal Collection Vehicle (SCV): (a) Schematic diagram of the internal structure of the perforating gun employed in this work; (b) Functional structure diagram of the SCV; (c) A SCV circuit board.
As shown in Fig. 2, the Signal Collection Vehicle (SCV) is designed for integrating into downhole tools. The SCV comprises several submodules, including the analog frontend (AFE) module, the signal processing and control module, the data storage module, the input-output (I/O) module, and auxiliary modules. Once the SCV begins operation, the AFE module samples CCL signals at 1 kHz via an analog-to-digital converter (ADC), converting them to digital format, and the control module stores these digital data in the storage module. Following completion of downhole operations, the SCV is salvaged to export data via the I/O module.
By sampling CCL signals downhole and eliminating long cable transmissions, signal integrity is preserved. The signals sampled by the SCV are designated as “original CCL waveforms” and form the foundation for dataset construction, as illustrated in Fig. 3(a).
Figure 3: The train and inference process of this work: (a) Segment normalized CCL waveform with manually labeled collar marks, each segment containing a collar mark at its center (signal shapes in boxes are illustrative only); (b) Augmentation methods for preprocessing waveform segments; (c)-(d) Neural network architectures employed in this work; (e) Multiple random augmentations applied to each segment generate diverse sub-segment variants for training and testing datasets; (f) The procedure of casing collar recognition from CCL waveforms using sliding windows with overlap.
II-C Dataset and Augmentation
Based on the original CCL waveforms, collar marks are manually annotated through expert analysis. For model training, CCL waveforms and their corresponding labels (in one-hot encoding) are segmented, with each segment containing a collar mark at its center, as illustrated in Fig. 3(a). Portions distant from collar marks are excluded.
@@ II-C Dataset and Augmentation
To enhance training effectiveness and maximize data utilization, we present various data augmentation methods for preprocessing original CCL data. These methods encompass data normalization, regularization, transformation and multiple sampling, which can be applied independently or in combination, as illustrated in Fig. 3(b).
@@
After preprocessing, the augmented segments and labels constitute the datasets used for model training and evaluation.
@@ II-C3 Label Smoothing Regularization (LSR)
where $p(k)$ is the ground-truth distribution over the kth class; $p^{\prime}(k)$ is the training distribution over the kth class; $i$ denotes the correct class; $K$ is the total number of classes; and $\delta$ is the Kronecker delta function.
This method discourages overconfident predictions by preventing the largest logit from becoming disproportionately larger than others, thereby regularizing the model and enhancing its generalization capability [29]. For this work, with $K=2$, the training distribution simplifies to:
@@ II-C4 Geometric Transformations
  • Time Scaling: Original sub-segments are scaled along the time axis by random factors and subsequently resampled to match the ADC sampling rate. To compensate for frequency distortion introduced by resampling, the Whittaker-Shannon interpolation formula (i.e., sinc interpolation) is chosen.
  • Randomly Cropping and Translation: These transformations segment waveforms into sub-segments that match both the sliding window length and the neural network model’s input length.
  • Amplitude Jittering: Original sub-segments are multiplied by random gain factors to enhance the model’s robustness and generalization capability.
  • Noise Injection: Gaussian noise is added to original sub-segments to improve model robustness against noise.
@@ II-C5 Multiple Sampling
To maximize dataset utilization, each segment undergoes multiple random augmentations to generate diverse sub-segment variants, as illustrated in Fig. 3(e). This approach theoretically accelerates convergence and enhance augmentation effectiveness.
@@ II-D Neural Network
The proposed model, Thin AlexNet (TAN), is a time-series version of the classical AlexNet architecture. The input and output dimensions are reduced from 2D to 1D to adapt CCL signals. Since AlexNet has demonstrated success in image classification [31, 32] and possesses sufficient simplicity to clearly demonstrate the effects of data augmentation, the TAN model serves as an appropriate benchmark.
@@ III-A Evaluation Measures
The CCL waveforms used in experiments originate from actual oil wells in Sichuan Province, China, ensuring that results are objective and representative of real-world conditions.
III-B Training and Testing
The training set comprises three-quarters of the CCL signal waveforms, with the remaining quarter allocated to the test set. The models are trained with a batch size of 16 for 100 epochs utilizing cross-entropy loss (CE loss) with an adaptive moment estimation (Adam) optimizer, implemented in PyTorch. Key configurations (Cfgs.) are presented in Tables I and II. The training progress and the optimal model performance for each configuration are illustrated in Fig. 4. Additionally, a mild interference waveform and a moderate interference waveform are evaluated through model inference, post-processing, and statistical collar identification performance analysis. The results for precision, recall, and F1 score are tabulated in Tables I and II.
III-C Results and Analysis
Figure 4: Evaluation metrics of training progress under different configurations, including cross-entropy loss, F1 score, and area under the precision-recall curve (AUC-PR). The curve indexes correspond to configurations in Tables I and II. The meanings of abbreviations refer to Tables I and II. (a)-(c) Evaluation metrics for Group 1 configurations; (d)-(f) Evaluation metrics for Group 2 configurations; (g)-(i) Enlarged sections of (d)-(f); (j)-(l) Evaluation metrics for Group 3 configurations; (m)-(o) Enlarged sections of (j)-(l); (p)-(r) legends for (a)-(c), (d)-(i), and (j)-(o), respectively.

### v3

# Data-Augmented Deep Learning for Downhole Depth Sensing and Validation

@@ Abstract
Accurate downhole depth measurement is essential for oil and gas well operations, directly influencing reservoir contact, production efficiency, and operational safety. Collar correlation using a casing collar locator (CCL) is fundamental for precise depth calibration. While neural network has achieved significant progress in collar recognition, preprocessing methods for such applications remain underdeveloped. Moreover, the limited availability of real well data poses substantial challenges for training neural network models that require extensive datasets. This paper presents a system integrated into a downhole toolstring for CCL log acquisition to facilitate dataset construction. Comprehensive preprocessing methods for data augmentation are proposed, and their effectiveness is evaluated using baseline neural network models. Through systematic experimentation across diverse configurations, the contribution of each augmentation method is analyzed. Results demonstrate that standardization, label distribution smoothing (LDS), and random cropping are fundamental prerequisites for model training, while label smoothing regularization (LSR), time scaling, and multiple sampling significantly enhance model generalization capabilities. Incorporating the proposed augmentation methods into the two baseline models results in maximum F1 score improvements of 0.027 and 0.024 for the TAN and MAN models, respectively. Furthermore, applying these techniques yields F1 score gains of up to 0.045 for the TAN model and 0.057 for the MAN model compared to prior studies. Performance evaluation on real CCL waveforms confirms the effectiveness and practical applicability of our approach. This work addresses the existing gaps in data augmentation methodologies for training casing collar recognition models under CCL data-limited conditions, and provides a technical foundation for the future automation of downhole operations.
@@ I Introduction
Accurately positioning downhole toolstrings (including perforating guns, bridge plugs, and packers) is essential in modern oil and gas well operations, directly affecting maximum productivity and operational safety [9]. Central to this task is precise downhole depth measurement, a challenge compounded by the extreme geometries of wellbores, as shown in Fig. 1 organized from [32], which often span thousands of meters in length while maintaining diameters of only a few inches [25].
Figure 1: Cross-sectional illustration of a typical oil and gas well structure. Representative casing collar signatures from magnetic response are illustrated in dark blue near the corresponding casing collar, while typical interference signals are illustrated in dark green. Organized from [32].
While surface wheel measurement (SWM) method offers a cost-effective means of estimating depth via a depth measuring head (DMH) as the toolstring descends during wireline intervention operations, it is susceptible to errors induced by cable slippage and elastic stretch. Furthermore, SWM is inapplicable to emerging operations such as wireless perforating. Consequently, to achieve accurate depth measurement in the absence of a DMH, depth correction using a casing collar locator (CCL, a magnetic downhole positioning tool that detects magnetic anomalies at the casing collars of wellbores or pipes) is fundamental [1]. The CCL produces a characteristic magnetic response as an electrical signal when passing through each collar [1, 8, 17], termed a “CCL response” or “collar (magnetic) signature”. The characteristic magnetic response pattern typically exhibits a bipolar signature, as illustrated by dark blue waveforms in Fig. 1. Through collar correlation, which refers to tying-in CCL logs with casing tally, depth reference markers are established, enabling actual and accurate depth measurement [19, 17]. The casing tally, also known as the list of collars or casing string reference depths, records the depths of casing collars and is commonly extracted from cementing quality data.
While collar correlation using CCL logs is an established method, collar signature recognition presents significant challenges. CCL signal integrity can be severely compromised by multiple factors, including cable effects, wellbore conditions [5], toolstring motion (swing or rotation), amplifier saturation, and environmental noise [19, 30]. Consequently, collar signature waveforms become increasingly ambiguous, necessitating robust recognition of collar signatures amid various interferences, as illustrated by dark green waveforms in Fig. 1.
@@
The limited sample amount and unique characteristics of CCL log data necessitate specialized preprocessing methods for neural network training. However, research on preprocessing methods remain scarce, despite extensive work on collar signature recognition [1, 30, 18, 16, 33, 11, 31, 21, 22]. Fortunately, extensive work exists on data augmentation for preventing overfitting and improving generalizability when training on small datasets. Data normalization methods – including min-max scaling, Z-score normalization (standardization), and robust scaling – demonstrate important roles in stabilizing input distributions and gradients, preventing vanishing and exploding gradients, and ultimately accelerating training convergence [10, 24, 3, 2]. Label smoothing regularization (LSR) and label distribution smoothing (LDS) discourage overconfident predictions, thereby enhancing model generalizability [34, 27]. Similarly, employing probability maps for boundary prediction instead of one-hot encoding (OHE) labels through boundary probabilization transforms the training objective from single-point prediction to probability distribution estimation, enabling models learn smooth and fuzzy decision boundaries that are more resistant to noise or perturbations [34, 23, 26]. Notably, Gaussian kernel often deliver optimal results in LDS applications [34]. Furthermore, graph augmentation techniques including randomly cropping, scaling, translation, and noise injection expand training datasets while improving model generalizability and robustness [12, 13].
@@
  • We develop a system integrated into downhole toolstring, called the Signal Collecting Vessel (SCV), illustrated in Fig. 2. The SCV samples raw CCL signals downhole and converts them to digital format, and stores them as waveforms for dataset construction.
  • We propose two neural networks models for collar signature recognition that serve as baselines for evaluating data preprocessing methods, as illustrated in Fig. 3. The first model, Thin AlexNet (TAN), is modified from AlexNet – a classic and proven architecture in pattern recognition. The second model, Miniaturized AlexNet (MAN), is a simplified version of TAN with fewer layers.
@@
  • We conduct extensive experiments across various configuration combinations with filed CCL logs to validate our methods. Results demonstrate that standardization, LDS, and random cropping are fundamental requirements for models training, while LSR, time scaling, and multiple sampling significantly enhance model generalization capability.
@@ II-A Problem Transformation
As previously discussed, accurate downhole toolstring positioning via collar correlation relies on the correct identification of bipolar patterns (i.e., collar signatures) within CCL logs. The centroid of each bipolar pattern is typically designated as the instant the CCL coincides with a casing collar. In the absence of DMH assistance (specifically, without the depth indexing commonly utilized in wireline logging), time-series CCL logging becomes the requisite approach.
Field practice demonstrates that collar signatures are identifiable primarily through local waveform characteristics in the vicinity of the collar, rendering distant signal features negligible. Consequently, analyzing CCL log fragments via an appropriately sized sliding window is an effective strategy, aligning with methodologies proposed in [23] and [14]. Furthermore, when raw CCL signals are sampled at a fixed frequency, absolute timestamps within the sliding windows become redundant; that is, the raw waveform sequence alone conveys sufficient information.
As shown in Fig. 1, the casing tally, which provides a series of depths, correlates to the centroids of collar signatures (hereafter referred to as “collar marks”) on the temporal axis. The most direct representation of collar marks in CCL waveforms is one-hot encoding (OHE), where a value of 1 indicates the presence of a collar mark and 0 indicates its absence. This formulation frames collar mark prediction as a binary classification task.
@@
However, collar marks exhibit extreme sparsity: Background samples (0s) outnumber target samples (1s) by several orders of magnitude, resulting in severe class imbalance. In backpropagation neural networks (BPNNs), this imbalance leads to sparse gradients and uneven penalty distribution, causing training instability and slow convergence. Consequently, training an effective classifier becomes computationally challenging. To mitigate this, OHE is replaced by a probability map [7, 23], wherein labels represent the probability of boundary occurrence. Interestingly, consistent with the theory in [7], the label reflects the distribution of relative importance across possible categories (i.e., “the presence of a collar mark”). This transformation shifts the problem from hard binary classification to boundary membership estimation, providing denser feedback during backpropagation and facilitating stable and efficient network training.
In summary, the problem of collar signature recognition is transformed into a boundary membership estimation task, where the input is a windowed temporal CCL waveform and the output is a temporal probability map.
II-B Acquisition of Raw CCL Waveforms
@@
Figure 2: Structure of downhole toolstring integrated with the Signal Collection Vehicle (SCV): (a) Schematic diagram of the internal structure of the perforating gun employed in this work; (b) Functional structure diagram of the SCV; (c) A SCV circuit board.
To eliminate signal degradation associated with long cable transmission and preserve signal integrity, a Signal Collection Vehicle (SCV) was developed, as shown in Fig. 2. The SCV is designed to be integrated into downhole toolstring and enable the real-time logging of raw CCL signals downhole.
The SCV comprises of the analog frontend (AFE) module, the signal processing and control module, the data storage module, the input-output (I/O) module, and auxiliary modules. As the SCV is lowered with downhole toolstring, the AFE module samples raw CCL signal at a sampling rate of $1\text{\,}\mathrm{kHz}$ via a 16-bit resolution analog-to-digital converter (ADC). The resulting digital data is recorded in the storage module. Upon completion of downhole operations, the SCV is salvaged, and data is exported via the I/O module.
The raw temporal CCL logs acquired by the SCV are designated as “original CCL waveforms”. Typically, a single log contains approximately 50 to 200 collar signatures, corresponding to a well depth ranging from $500\text{\,}\mathrm{m}$ to $2\text{\,}\mathrm{km}$.
II-C Dataset Construction and Augmentation
Collar marks within the field-acquired original CCL waveforms were manually annotated through expert analysis. To construct the dataset, waveforms surrounding each collar mark were fragmented into fixed-length fragments, while non-informative sections distant from the collars were excluded. As illustrated in Fig. 3(a), each fragment is centered on a collar mark, which is initially represented using one-hot encoding.
@@ II-C Dataset Construction and Augmentation
To enhance training effectiveness, maximize data utilization, and avoid overfitting, we present various data augmentation methods for preprocessing original CCL data. These methods encompass data normalization, regularization, transformation and multiple sampling, which can be applied independently or in combination, as illustrated in Fig. 3(b).
@@
After preprocessing, the augmented waveform fragments and target labels constitute the datasets used for model training and evaluation.
@@ II-C3 Label Smoothing Regularization (LSR)
where $p(k)$ is the ground-truth distribution over the kth class; $p^{\prime}(k)$ is the training distribution over the kth class; $i$ denotes the correct class; $K$ is the total number of classes; $\delta$ is the Kronecker delta function; and $\epsilon$ is the smoothing parameter.
When employing the sigmoid function, the parameter $\epsilon$ effectively constrains the magnitude of the output logits. This mechanism discourages overconfident predictions by preventing the largest logit from becoming disproportionately larger than others, thereby regularizing the model and enhancing its generalization capability [27]. In this study, given $K=2$, the training distribution simplifies to:
@@
where an $\epsilon$ value of 0.1 is recommended to strike a balance between accuracy and generalization capability.
@@ II-C4 Geometric Transformations
Figure 3: The train and inference process of this work: (a) Fragment the normalized CCL waveform based on manually labeled collar marks, each fragment containing a collar mark at its center (waveform shapes in boxes are illustrative only); (b) Augmentation methods for preprocessing waveform fragments and their labels; (c)–(d) Baseline neural network architectures employed in this work; (e) Multiple random augmentations applied to each fragment to generate diverse sub-samples variants for training and testing datasets; (f) The procedure of casing collar recognition from CCL waveforms using sliding windows with overlap.
@@
  • Time Scaling: Waveform fragments are scaled along the time axis by random factors and subsequently resampled to restore the original sampling rate. The resampling process employs Hann-windowed sinc interpolation, the default resampling method in the TorchAudio library, to mitigate spectral artifacts, including ringing and aliasing, while ensuring effective high-frequency attenuation.
  • Randomly Cropping and Translation: These transformations waveform fragments into sub-samples that match both the sliding window length and the neural network model’s input length.
  • Amplitude Jittering: Waveform fragments are multiplied by random gain factors to enhance the model’s robustness and generalization capability.
  • Noise Injection: Gaussian noise is added to original fragments to improve model robustness against noise.
@@ II-C5 Multiple Sampling
To maximize dataset utilization, each fragments undergoes multiple random augmentation process to generate diverse sub-samples variants, as illustrated in Fig. 3(e). This approach theoretically accelerates convergence and enhance augmentation effectiveness.
@@ II-D Neural Network
The proposed model, Thin AlexNet (TAN), is a time-series version of the classical AlexNet architecture. The input and output dimensions are reduced from 2D to 1D to adapt CCL waveforms. Since AlexNet has demonstrated success in image classification [12, 13] and possesses sufficient simplicity to clearly demonstrate the effects of data augmentation, the TAN model serves as an appropriate baseline.
@@ III-A Evaluation Measures
III-B Training and Validation
The CCL waveforms utilized in the experiments were acquired from field operations in Sichuan Province, China, ensuring that the results are representative of actual downhole conditions. The dataset was partitioned into training and validation subsets using a 3:1 ratio based on the CCL logs, yielding 288 and 50 original waveform fragments, respectively. To mitigate the limitations (including funds) imposed by the restricted dataset size, a multiple sampling strategy was employed to augment the training set. The training and validation processes were conducted offline on a workstation using the PyTorch framework.
Figure 4: Evaluation metrics of training progress under different configurations, including cross-entropy loss, F1 score, and area under the precision-recall curve (AUC-PR). The curve indexes correspond to configurations in Tables I and II. The meanings of abbreviations refer to Tables I and II. (a)–(c) Evaluation metrics for Group 1 configurations; (d)–(f) Evaluation metrics for Group 2 configurations; (g)–(i) Enlarged sections of (d)–(f); (j)–(l) Evaluation metrics for Group 3 configurations; (m)–(o) Enlarged sections of (j)–(l); (p)–(r) legends for (a)–(c), (d)–(i), and (j)–(o), respectively.